OpenStudy (anonymous):

Doug walks at a steady pace of 5280 feet per hour. While Tommy walks at a steady pace of 3/4 of that of Doug. After one half hour, how far apart are they?

OpenStudy (anonymous):

I dont know the equation to solve it. and I am very bad at solving fractions :(

OpenStudy (magbak):

2 miles.

OpenStudy (anonymous):

thanks I was kind of hoping to see how to solve it too

OpenStudy (espex):

The distance (d) between the two is growing at a rate of doug (\(\frac{5280~feet}{1~hour}\)) plus the rate of Tommy (\(\frac{3}{4}*doug\)). To see how far apart they get, take the distance and multiply it by the time in hours that they walk (\(\frac{1}{2}hour*d\)). Substituting and expanding you get:\[\frac{1~hour}{2}*d \rightarrow \frac{1~hour}{2}*(\frac{5280~feet}{1~hour}+ (\frac{3~hour}{4}*doug))\]\[\rightarrow \frac{1~hour}{2}*(\frac{5280~feet}{1~hour}+ (\frac{3~hour}{4}*\frac{5280~feet}{1~hour}))\]\[\rightarrow\frac{1~hour}{2}*(\frac{5280~feet}{1~hour}+\frac{15840~hour~feet}{4~hour})\rightarrow \frac{1~hour}{2}*(\frac{5280~feet}{1~hour}+\frac{3960~feet}{1})\]\[\rightarrow\frac{1~hour}{2}*\frac{9240~feet}{1~hour} \rightarrow 4620~feet\] Given a conversion factor of \(\frac{1~mile}{5280~feet}\), you end up with \(\frac{4620}{5280}~mile\rightarrow 0.875~miles\)